Ahat is the Study of alectno Caridee Cardiovascular clisease is the of seath in America Cause hat are the unchangeale risk facrs re arding Cordis tas cular diaease galtanometer G Name the sclentisk that euop e nente are toaus b calculade the heact tats
Answer #1

1. what is the study of electrocardiology?

Electrocardiology is the study of the electrical signals of the heart through the electrodes. These signals show the rhythm of the heart. It helps to determine the diagnostic condition of the heart and to provide appropriate treatment.

2. Cardiovascular disease is the leading cause of death in America.

In the U.S Around 45.1% death are due to Coronary Heart disease, 16.5% are due to stroke, 9.1% death are due to high blood pressure, and the 3.2% account for artery diseases and the 8.5% death are due to heart failure.

3. What are the unchangeable risk factors regarding Cardiovascular disease?

  • Age - Elderly people more likely to get heart disease than young.
  • Gender - Men are the greater risk of getting heart disease than women.
  • Family History - Those who have the family history of heart disease in parents or close relatives are more prone to get it.

4. What are the reasons to perform EKG?

  • Chest pain
  • Pain in the neck, arm, back.
  • Syncope
  • Cyanosis
  • Epigastric pain
  • Diaphoresis
  • External dyspnea

5. Name the scientist who developed galvanometer?

Galvanometer was developed by Hans Oersted in 1820. Later Marcel Deprez and Dr.Arsonval developed the mirror type to measure the electric current. Electromagnetic Galvanometer was developed by Johann Schweigger in 1820. Then the original mirror type was developed by Johann Christian in 1826.

6. What are the layers of the heart?

  • Epicardium - the outermost layer
  • Myocardium - middle layer
  • Endocardium - inner layer

7. What are the different names of Limb leads?

Bipolar Limb lead

  • Lead I RA to LA
  • Lead II RA to LL
  • Lead III LA to LL

Unipolar Limb Lead

  • Lead aVR RA to LA and LL
  • Lead aVL LA to RA and LL
  • Lead aVF LL to RA and LA.

8. What are the ways to calculate heart rate?

The best way to calculate heart rate from EKG is

1. Count the number of R-R interval in 6 seconds rhythm strip and then multiply by 10.

Ex: If the total of R-R interval is 7, then 7x 10 = 70bpm.

2. Count the number of big boxes between the R-R interval. One big box is 0.2 sec. Then 60 will be divided by the number of big boxes. Rate = 60/ R-R interval

For Ex:

1 big box = 60/0.2 = 300bpm

2 big box = 60/0.4 = 150bpm

3 big box = 60/0.6 = 100bpm

4 big box = 60/0.8 = 75bpm